OpenStudy (anonymous):

A local animal shelter currently has 18 dogs and 24 cats available for adoption. Six of the dogs are male. If the species and sex of the animals are independent, how many cats are female? A) 6 B) 8 C) 12 D) 16

OpenStudy (anonymous):

I don't even think this is possible to work out.

OpenStudy (carlyleukhardt):

C i think

OpenStudy (anonymous):

@FitzIncorporated @carlyleukhardt Why?

OpenStudy (anonymous):

Not enough information.

OpenStudy (anonymous):

@Rushwr What do you think?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

I also think it is C

OpenStudy (anonymous):

@lauren.ascough Can you please explain?

OpenStudy (rushwr):

I guess it's D

OpenStudy (anonymous):

Can someone please explain this answers?

OpenStudy (rushwr):

there are 18 dogs 6 of them are male and 12 are female so write it as a fraction of the total number of dogs. \[\frac{ 12 }{ 18 } = \frac{ 2 }{ 3 }\] So we can say 2/3 of each species is female right? if the total number of cats are 24 then 2/3 of it is female \[24 * \frac{ 2 }{ 3} = 16\] This is how I got the answer !

OpenStudy (anonymous):

Thank you!!